### v1

# From Light-Cone to Supersonic Propagation of Correlations by Competing Short- and Long-Range Couplings

Abstract
We investigate the dynamical spreading of correlations in many-body quantum systems with competing short- and global-range couplings. We monitor the non-equilibrium dynamics of the correlations following a quench, showing that for strong short-range couplings the propagation of correlations is dominated at short and intermediate distances by a causal, light-cone, dynamics, resembling the purely short-range quantum systems. However, the interplay of short- and global-range couplings leads to a crossover between space-time regions in which the light-cone persists to regions where a supersonic, distance-independent, spreading of the correlations occurs. We identify the important ingredients needed for capturing the supersonic spreading and demonstrate our findings in systems of interacting bosonic atoms, in which the global range coupling is realized by a coupling to a cavity light field, or atomic long-range interactions, respectively. We show that our results hold in both one and two dimensions and in the presence of dissipation. Furthermore, we characterize the short time power-law scaling of the distance-independent growth of the density-density correlations.
A fundamental question attracting enormous interest in recent decades is whether there are limits on how fast physical effects can propagate in quantum systems Lieb and Robinson (1972); Chen et al. (2023); Defenu et al. (2024, 2023). Beside its fundamental character, the answer to this question is also relevant for quantum technologies, in particular, for the preparation time of sought-after states, quantum information links and for quantum operations more generally. In this context obtaining deep insights into how correlations spread on different platforms enables one to select the most suitable one based on their speed and versatility in performing desired applications.
In typical quantum systems with short-range interactions the seminal work of Lieb and Robinson has shown that correlations can only travel with a velocity that is bounded from above Lieb and Robinson (1972). The Lieb-Robinson bound can be applied to various topics Chen et al. (2023), e.g. the area law of entanglement Hastings (2007); Van Acoleyen et al. (2013), the exponential clustering for correlation functions Nachtergaele and Sims (2006), or the scrambling of information Roberts and Swingle (2016). Although the original derivation of this bound is not strictly applicable to bosonic systems, an effective light cone also exists for them at low densities Takasu et al. (2020); Kuwahara and Saito (2021a); Yin and Lucas (2022) and in the Bose-Hubbard model Kollath et al. (2007); Läuchli and Kollath (2008); Cheneau et al. (2012); Barmettler et al. (2012); Carleo et al. (2014); Faupin et al. (2022a, b). Beyond light-cone propagation is typically only present in tailored models  Eisert and Gross (2009); Kuwahara et al. (2024); Wang and Hazzard (2020). The typical light-cone spreading of correlations for short-range couplings has been observed experimentally with ultra-cold atoms in optical lattices Cheneau et al. (2012); Barmettler et al. (2012).
The robustness of the light-cone spreading of correlations in systems with short-range interactions opened the question whether long-range interactions lead to a different propagation of the correlations, since they directly connect distant sites Defenu et al. (2024, 2023). An important class of long-range interactions are algebraically decaying, i.e. $x^{-\alpha}$, present in trapped ions, Rydberg, and dipolar systems Defenu et al. (2023). The role of algebraically decaying interactions for correlation spreading has been experimentally investigated with trapped ions Richerme et al. (2014); Jurcevic et al. (2014). While, in general, the concept of Lieb-Robinson bounds is not applicable to arbitrary long-range interacting systems Hastings (2012); Eisert et al. (2013); Foss-Feig et al. (2015); Schneider et al. (2021); Vu et al. (2024), other approaches pointed out that for sufficiently large values of $\alpha$ Eisert et al. (2013); Hauke and Tagliacozzo (2013); Cevolani et al. (2015); Chen and Lucas (2019); Tran et al. (2020); Kuwahara and Saito (2020); Foss-Feig et al. (2015); Else et al. (2020); Chen and Lucas (2021); Tran et al. (2021); Kuwahara and Saito (2021b); Faupin et al. (2022a, b); Gong et al. (2023); Vu et al. (2024), or special setups Jünemann et al. (2013), the correlation propagation is still effectively bound.
Still, experimentally relevant examples where unbounded correlation spreading is observed are rare and the transition from light-cone like dynamics to the latter by competing short- and long-range interaction is not yet explored. For instance, an open question in this field is whether in the presence of long-range interactions, with sufficiently small $\alpha$, their effect will dominate the shorter-range processes and simply wash out all notion of locality leading to an instantaneous build up of correlations over all distances.
In this work, we investigate the emergence and disappearance of the light-cone dynamics in the presence of competing short- and global-range couplings in an out-of-equilibrium setting. We show that light-cone dynamics can trigger the onset of non-causal propagation of the globally interacting dynamics, where correlations can spread across the system almost simultaneously, independent of the distance. We dub this regime supersonic, distance-independent, propagation. As in such scenarios a distance-dependent bound is absent due to the long-range interactions we investigate the dynamics of correlations by performing numerically exact simulations for systems with short- and global-range couplings, in the presence and absence of dissipation. We find that a crossover between a light-cone behavior and the supersonic spreading of correlations arises. We supplement our simulations based on matrix product states with approximate numerical and analytical techniques to understand the key ingredients necessary for the supersonic propagation. We identify that the fluctuations of the global coupling act as the carriers of the long-range correlations, however, the exact nature of these fluctuations is not important. In particular, we show that the supersonic distance-independent propagation appears for a Bose-Hubbard atomic model, in the presence of: (i) a global-range coupling to a bosonic mode; (ii) a global purely atomic interaction; (iii) a global fluctuating field. Additionally, our findings are robust against dissipation and independent from the dimensionality of the system, making the observed effects experimentally realizable, allowing one to explore the interplay of correlations spreading and non-local dissipation.
Setup: We consider one of the most prominent platforms for controlling the interplay between short- and long-range couplings, interacting bosonic atoms confined to an external optical lattice and coupled to the field of a high-finesse cavity Ritsch et al. (2013); Mivehvar et al. (2021). A lot of theoretical and experimental effort has been focused on the nature of steady states in such models Klinder et al. (2015); Landig et al. (2016); Hruby et al. (2018); Ritsch et al. (2013); Niedenzu et al. (2010); Silver et al. (2010); Fernández-Vidal et al. (2010); Li et al. (2013); Bakhtiari et al. (2015); Flottat et al. (2017); Lin et al. (2019); Himbert et al. (2019); Halati et al. (2020a); Bezvershenko et al. (2021); Sharma et al. (2022); Chanda et al. (2022), and certain aspects of their dynamics Rodríguez Chiacchio and Nunnenkamp (2018); Halati et al. (2022, 2025). The evolution of the density matrix $\hat{\rho}$ is described by the Lindblad master equation Carmichael (1991); Breuer and Petruccione (2002); Maschler et al. (2008); Ritsch et al. (2013); Mivehvar et al. (2021)
$\partialderivative{t}\hat{\rho}=-\frac{i}{\hbar}\left[\hat{H},\hat{\rho}\right]+\frac{\Gamma}{2}\left(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\hat{a}^{\dagger}\hat{a}\hat{\rho}-\hat{\rho}\hat{a}^{\dagger}\hat{a}\right),$ (1)
where $\Gamma$ is the strength of the cavity losses. The Hamiltonian is given by
$\hat{H}=\hat{H}_{\text{cav}}+\hat{H}_{\text{BH}},~{}\hat{H}_{\text{cav}}=\hbar\delta\hat{a}^{\dagger}\hat{a}-\hbar\Omega(\hat{a}+\hat{a}^{\dagger})\hat{\Delta},$ (2)
[…]
The local processes consist in the repulsive on-site interactions of strength $U$ and atomic tunneling with the amplitude $J$ between neighboring sites $\langle j,j^{\prime}\rangle$. The effective light-matter coupling strength $\Omega$ is controllable by a transverse standing-wave laser beam, where $\delta$ is the detuning of the cavity with respect to the pump frequency. The period of the lattice is chosen to be twice the period of the cavity mode. This effectively couples the cavity to the density imbalance of the atoms of a bipartite lattice, $\hat{\Delta}=\sum_{j\in A}\hat{n}_{j}-\sum_{j\in B}\hat{n}_{j}$, with the sublattices $A$ and $B$.
Figure 1: (a), (c), (e) The space-time propagation of the correlations $\mathcal{C}_{nn}(d,t)$, computed with $\hat{H}$, Eq. (2), for the atoms-cavity coupling strength $\hbar\Omega/J\in\{0.03,0.42,0.99\}$. The blue line is a guide to the eye and approximates the front of the light-cone propagation for $\Omega=0$. (b), (d), (f) The dynamics of $\mathcal{C}_{nn}(d,t)$ for several distances and $\hbar\Omega/J\in\{0.03,0.42,0.99\}$. The curves depicted with shades of orange to brown correspond to $\hat{H}$, Eq. (2), the dashed curves of shades of green to $\hat{H}_{\text{atom-only}}$, Eq. (4), and the blue curves to the $\Omega=0$ case. The magenta lines represent algebraically scaling $\propto t^{\alpha}$, with $\alpha\in\{2,4,8\}$. The parameters used are $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$, $\hbar\Gamma/J=0$.
Quench scenario: We simulate the propagation of the correlations in a quench scenario in which initially the atoms are in an uncorrelated product state with one atom every two sites, |$\ket{1010\dots}$, and the cavity is empty. The initial state breaks the $\mathbb{Z}_{2}$ symmetry of $\hat{H}$, $(\hat{a},\hat{\Delta})\!\to\!(-\hat{a},-\hat{\Delta})$. We analyze the dynamics of connected density-density correlations
$\mathcal{C}_{nn}(d,t)=\frac{1}{\mathcal{N}}\sum_{j,j^{\prime},|j-j^{\prime}|=d}\left|\left\langle\hat{n}_{j}\hat{n}_{j^{\prime}}\right\rangle-\left\langle\hat{n}_{j}\right\rangle\left\langle\hat{n}_{j^{\prime}}\right\rangle\right|(t),$ (3)
where we average over all the correlations at a certain distance $d\equiv|j-j^{\prime}|$ with $\mathcal{N}$ being the number of sites at the given distance $d$. In $\mathcal{C}_{nn}(d,t)$ we subtract the disconnected part of the density-density correlations, $\left\langle\hat{n}_{j}\right\rangle\left\langle\hat{n}_{j^{\prime}}\right\rangle$, which describes the reorganization of the atomic density in the cavity-induced potential.
As in the initial state the different sites are uncorrelated, this scenario allows us to investigate cleanly the space-time propagation of correlations through the system. Let us remark that one particularity of the considered correlations is that in the limit of vanishing tunneling, $J\!\!=0$, $\hat{H}$ commutes with density operators and the correlations $\mathcal{C}_{nn}$ would be constant in time. We investigate the propagation of correlations for $J\!\neq\!0$, in the presence of competing short-range and global-range terms for the 1D version of $\hat{H}$, Eq. (2), and of the Liouvillian, Eq. (1), using a recently developed method based on time-dependent matrix product state (tMPS) techniques employing swap gates for the cavity coupling and the quantum trajectories approach for the dissipative dynamics Halati et al. (2020b); supp .
Light-cone to supersonic evolution: We now investigate the crossover between the light-cone evolution and distance-independent propagation of the correlations in Fig. 1, for the 1D chain of atoms coupled to the optical cavity, assuming no cavity losses ($\Gamma\!=\!0$). For low coupling to the cavity field [Fig. 1(a) for $\hbar\Omega/J=0.03$] and the considered distances a light-cone propagation is found following the blue line, which marks approximately the speed limit for the propagation of the correlations we found in the absence of the global-range coupling. For the same value of the coupling, we present in Fig. 1(b) $\mathcal{C}_{nn}(d,t)$ for certain distances as a function of time. The correlations exhibit an algebraic increase with time until reaching a maximum following the light-cone. By comparing with the $\Omega=0$ case we observe for correlations at longer distances, $d\gtrsim 10$, finite values outside of the light-cone. This can be understood by expanding the time-evolution operator, i.e. $e^{-i\hat{H}t/\hbar}\!\approx\!e^{-i\hat{H}_{\text{BH}}t}+O(t^{2}\Omega)$. Thus, the time-evolution is dominated by the Bose-Hubbard evolution beside corrections on the order of $O(t^{2}\Omega)$, or higher depending on the initial state and the form of the coupling supp .
The competing behavior becomes much more apparent as we increase the strength of the coupling to the cavity, $\hbar\Omega/J=0.42$ in Fig. 1(c)-(d). In Fig. 1(c) we observe a coexistence of the light-cone induced by short-range interaction and a distance-independent increase of correlations, for $d>4$, due to the presence of the cavity-induced global-range interactions. To be more precise for the time-evolution of the correlations at different distances in Fig. 1(d) we have an initial increase of $\mathcal{C}_{nn}(d,t)$ for $d\leq 3$ one after the other, while for $d\geq 5$ all correlations increase simultaneously. The fact that we observe at very short distances and times $tJ/\hbar\lesssim 1$ the emergence of correlations following the light-cone is due to our choice of the initial state and observables. The density correlations commute with the entire Hamiltonian except the tunneling terms, implying that initially a change in the correlations can only be induced by tunneling. Thus, the cavity-mediated long-range interactions can start spreading the correlations only triggered by the short-range hopping term. By increasing the coupling strength to the cavity even further, $\hbar\Omega/J\!=\!0.99$ in Fig. 1(e)-(f), we observe that the instantaneous spreading of the correlations occurs at a shorter time, compared to the parameters in Fig. 1(c)-(d), and only for $d=1$ and $d=2$ the correlations start growing beforehand. Furthermore, for such strong long-range interactions we do not observe any remnant of the light-cone dynamics, the role of the short-range processes being restricted to generating locally the density-density correlations, which propagate via the cavity coupling.
Figure 2: (a)-(b) The dependence of the correlations $\mathcal{C}_{nn}$ as a function of rescaled time $t\sqrt{\Omega J/\hbar}$ for (a) $d=5$ and (b) $d=16$, obtained for the atom-cavity Hamiltonian $\hat{H}$, Eq. (2). (c)-(d) The dependence of the correlations $\mathcal{C}_{nn}$ as a function of rescaled time $t\Omega^{2/3}(J/\hbar)^{1/3}$ for (a) $d=5$ and (b) $d=16$, obtained for the atom-only Hamiltonian $\hat{H}_{\text{atom-only}}$, Eq. (4). The magenta lines represent the algebraic scaling (a)-(b) $\propto t^{8}\Omega^{4}$ and (a)-(b) $\propto t^{6}\Omega^{4}$. The values of the coupling are $0.03\leq\hbar\Omega/J\leq 1.02$, and the same parameters as in Fig. 1.
The minimal ingredients necessary to understand the distance-independent spreading are a simplified tunneling term and the global coupling to the cavity supp . Considering $\hbar\Omega\gg J$ the correlations build up following the scaling $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}J^{4}t^{8}\left[\langle(\hat{a}+\hat{a}^{\dagger})^{2}\rangle-\langle\hat{a}+\hat{a}^{\dagger}\rangle^{2}\right]$, if the initial state is an uncorrelated Fock state, where the contribution is proportional to the fluctuations present in the cavity field. In Fig. 2(a),(b) we show that this scaling is dominant in the large coupling regime taking the atoms-cavity model $\hat{H}$ by collapsing the numerical results for different values of the coupling $\Omega$ onto a single curve using $(t\sqrt{\Omega J})^{8}$. Let us note that this scaling changes when other energy scales, such as the full kinetic term, the detuning $\delta$, or preexisting long-range density correlations in the initial state, become relevant.
Generalization to atom-only global interactions: In order to be able to make a more general statement regarding the phenomenology of the spreading of correlations in systems with both short-range and global-range interactions, we contrast the results of the coupled atoms-cavity system with the following atom-only model,
$\hat{H}_{\text{atom-only}}=\hat{H}_{\text{glo}}+\hat{H}_{\text{BH}},~{}\hat{H}_{\text{glo}}=-\frac{\hbar\Omega^{2}\delta}{\delta^{2}+\Gamma^{2}/4}\hat{\Delta}^{2}.$ (4)
This model represents an atom-only description of the complex hybrid system obtained by eliminating the cavity field in the limit $J,U,\hbar\Omega^{2}/\delta\ll\hbar\delta$ Ritsch et al. (2013); Mivehvar et al. (2021), and can include global range dissipative processes Fernández-Vidal et al. (2010); Mivehvar et al. (2021); Rodríguez Chiacchio and Nunnenkamp (2018); Jäger et al. (2022). $\hat{H}_{\text{atom-only}}$ includes directly the effective global-range nature induced by the coupling to the cavity in $\hat{H}_{\text{glo}}$. In the following, we complement the results obtained with the full atoms-cavity Hamiltonian, $\hat{H}$, with results obtained with $\hat{H}_{\text{atom-only}}$. The simulations of the atom-only model, Eq. (4), make use of the matrix product states implementation of the time-dependent variational principal (TDVP) Haegeman et al. (2011, 2016); supp . We show that even if the global range coupling has a different form, stemming from purely atomic global interactions, the same crossover in the correlation spreading occurs as for the coupling to the cavity field.
In Fig. 1(b),(d), showing the dynamics of $\mathcal{C}_{nn}(d,t)$ for several distances, we plot both the results for the atoms-cavity Hamiltonian, Eq. (2), (continuous lines), and the results for the corresponding atom-only model, Eq. (4), (dashed lines). At small values of the long-range interactions, Fig. 1(b), we obtain a very good quantitative agreement in the dynamics of the density-density correlations. In Fig. 1(d) we observe that the initial rise of the correlations $d=1$ and $d=2$ agrees in the two models, confirming that they are mostly influenced by the short-range terms. While in both cases we have a simultaneous increase of $\mathcal{C}_{nn}(d>3,t)$, the rise of $\mathcal{C}_{nn}(d>3,t)$ for the atom-only model occurs at slightly earlier times than for the atoms-cavity model. In particular, the scaling we find is following more closely $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}t^{6}$, as shown in Fig. 2(c),(d). We note that by decreasing the photonic time-scale, i.e. increasing $\delta$, while keeping $\Omega^{2}/\delta$ constant, the difference between the long distance correlations of the two models decreases (not shown) and eventually they agree in the limit in which $\hbar\delta$ is the largest energy scale. In this regime also the atoms-cavity model shows the scaling $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}t^{6}$, implying that the main effect of the fast photons is to introduce the effective global interaction. The qualitative agreement in the behavior of the correlations in the two models in this regime shows that the main features of the competition of the light-cone and the distance-independent propagation due to the global range interactions are more generic than the case of hybrid atoms-cavity systems.
Figure 3: (a) The space-time propagation of the correlations $\mathcal{C}_{nn}(d,t)$, in the presence of dissipation, Eqs. (1)-(2), for $\hbar\Omega/J=0.99$, $\hbar\Gamma/J=10$, $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$. The blue line is a guide to the eye and approximates the front of the light-cone propagation for $\Omega=0$ and $\Gamma=0$. (b) The propagation of correlations when cavity field is replaced with a classical field supp , for the same parameters as in panel (a). Correlations calculated from TWA as function of time and distance in (c) 1D for a lattice with L=21 and in (d) 2D for lattice with $L\times L=21\times 21$. For the TWA simulations we have used $U/J=0.1$, $\hbar\delta/J=2$, $\hbar\Gamma/J=1$, $\sqrt{N}\hbar\Omega/J=4$ in (c), and $\sqrt{N}\hbar\Omega/J=8$ in (d). For the simulations we initialize the atoms with alternating densities of $n=10$ and $n=11$ bosons corresponding to the total atom numbers (c) $N=220$ and (d) $N=4630$.
Influence of a dissipative cavity field: A seldom explored question is how the presence of dissipation alters the spreading of correlations. Seminal works showed that for local and quasi-local Lindblad operators Lieb-Robinson-like bounds exist Poulin (2010); Schuch et al. (2011); Barthel and Kliesch (2012); del Campo et al. (2013); Funo et al. (2019), observed these numerically Bernier et al. (2018) and characterized the dynamics of mutual information Alba and Carollo (2021) and operator entanglement Wellnitz et al. (2022). The dissipation range has been shown to be able to alter the spatio-temporal behavior of correlations in spin systems Seetharam et al. (2022). Thus, it is an interesting question to understand the effects of dissipation in form of cavity losses, Eq. (1), which is a highly non-local dissipation for the atoms due to the global coupling.
In the present work, the dissipation can be a key player in changing the position of crossover between the local and the global spreading of correlations. By increasing the strength of dissipation we can go from the regime in which the distance-independent propagation of correlations dominates to a regime with a mostly light-cone evolution [see Fig. 3(a) compared to Fig. 1(e)-(f)]. Thus, whereas the dissipation globally couples to the atoms, the photon losses rescale the effective global atom-atom interactions to a lower value [see the coefficient in $\hat{H}_{\text{glo}}$ Eq. (4)], making the light-cone propagation more prominent. However, for the same strength of $\hat{H}_{\text{glo}}$ stronger dissipation helps the supersonic propagation.
Key ingredients for the supersonic propagation: To highlight the required ingredients for competition between the light-cone and distance-independent spreading of the correlations, we show in Fig. 3(b) results obtained employing a mean-field description of the cavity field. Within this approach the cavity mode is modeled by a time-dependent classical field coupled to the mean value $\langle\hat{\Delta}\rangle$ supp . We observe that in this simplified approach we cannot recover the evolution shown in Fig. 3(a). Whereas the light-cone like evolution is evident also in Fig. 3(b), the distance-independent spreading of correlations is absent. As the cavity field couples solely to the mean value $\langle\hat{\Delta}\rangle$ it looses the ability to transport and create fluctuations. However, by including stochastic noise arising from photon losses in the evolution of the classical field we recover a distance-independent propagation, albeit with a reduced contribution than in Fig. 3(a) supp . Thus, the fluctuations of the cavity field are of crucial importance.
Spreading of correlations at higher fillings and two-dimensions: In order to obtain a simplified description which captures the ability of the cavity field to transport density fluctuations, we employ the truncated Wigner approximation (TWA). In the TWA the cavity field and each atomic site is described as a complex stochastic classical field Gardiner et al. (2002); supp and it has been previously shown to describe light-cone spreading of correlations in Bose-Hubbard models Nagao et al. (2019). This approach allows us to extend our tMPS results to regimes of high atomic occupancies and to higher dimensions.
For the 1D case at high filling, we show in Fig. 3(c) the TWA results, for an initial state with a staggered atomic occupation alternating between $n\!=\!10$ and $n\!=\!11$ supp . We highlight that the qualitative behavior is the same as discussed in Fig. 1(e) for the tMPS results, at short times light-cone spreading of the correlations occurs, followed by a distance-independent rise of correlations at large distances. Establishing the TWA as a valid approach to simulate the quantum dynamics of correlations in this model gives us the opportunity to study this effect in higher dimensions. In Fig. 3(d), we show the results for a 2D atomic system coupled to the cavity field, where the atomic initial state is a checkerboard with alternating $n\!=\!10$ and $n\!=\!11$ occupations. Compared to the 1D simulations the magnitude of $\mathcal{C}_{nn}$ is reduced, which we attribute to radial spread of the correlations in 2D. Nevertheless, the key features that emerge from the competing short- and long-range processes are well visible in 2D, the light-cone propagation of correlations on short times and the supersonic spreading on longer times.
Conclusions: We investigated the dynamics of density-density correlations emerging from the interplay of short- and global-range couplings, finding two key features, a light-cone propagation at short distances and times followed by a supersonic, distance-independent, spreading. The crossover time between these dynamical features and their relative strength depend crucially on the ratio between the global coupling and the short-range tunneling. Furthermore, the general character of our results is highlighted by the robustness to the different quench situations, the dimensionality of the system, the presence of dissipation, or the form of the global coupling. We identify that the emerging dynamics relies on the presence of fluctuations in the long-range couplings, stemming either from the cavity field or the interaction itself. The role of the fluctuations is essential for propagating the correlations outside of the light-cone, which further underline the importance of capturing the fluctuations in the atoms-cavity coupling in obtaining the correct quantum dynamics, as shown also in our previous works Halati et al. (2020a, 2022, 2025). The effect could be realized and investigated in current state-of-the-art experiments of ultra-cold atoms coupled to optical cavities. However, due to its generality our findings are applicable to a much wider class of experimental systems. We emphasize that our results cannot be inferred from considerations based on general bounds on the propagation of correlations, which would be dominated by the contributions of the global-range couplings, such that numerical simulations of the quantum dynamics were needed. Our results can guide the experimental investigations in controlling the spreading of correlations in platforms with competing interaction scales.
Acknowledgments: We thank J.P. Brantut and T. Donner for fruitful discussions. We acknowledge support by the Swiss National Science Foundation under Division II Grant No. 200020-219400 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 277625399-TRR 185 OSCAR (“Open System Control of Atomic and Photonic Matter”, B4), No. 277146847-CRC 1238 (“Control and dynamics of quantum materials”, C05), CRC 1639 NuMeriQS (“Numerical Methods for Dynamics and Structure Formation in Quantum Systems”) – project No. 511713970 , Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”), and under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769. This research was supported in part by the National Science Foundation under Grants No. NSF PHY-1748958 and PHY-2309135.
Data availability: The supporting data for this article are openly available at Zenodo datazenodo .
A Time-dependent matrix product states method for long-range quantum systems
The numerically exact simulations for the time evolution of the Liouvillian of a one-dimensional Bose-Hubbard model coupled to a dissipative cavity [Eqs. (1)-(2) in the main text] have been performed by employing a recent implementation of a matrix product states (MPS) method Halati et al. (2020a, b); Halati (2021). The time-evolution of the Hamiltonian which contains the global-range coupling to the cavity makes use of a variant of the quasi-exact time-dependent variational matrix product state (tMPS) based on the Trotter-Suzuki decomposition of the time evolution propagator White and Feiguin (2004); Daley et al. (2004); Schollwöck (2011) and the dynamical deformation of the MPS structure using swap gates Stoudenmire and White (2010); Schollwöck (2011); Wall et al. (2016). For the results in the presence of cavity losses we employ the stochastic unravelling of the master equation with quantum trajectories Dalibard et al. (1992); Gardiner et al. (1992); Daley (2014). The method has been implemented by employing the ITensor Library Fishman et al. (2022). Additional details regarding the implementation and benchmarks can be found in Ref. Halati et al. (2020b).
In order to ensure the convergence of the density-density correlations, $\mathcal{C}_{nn}$, [see main text Eq. (3) for their definition], up to an error of approximately ${}10^{-5}$ for the times considered in the results presented in the absence of dissipation, we chose the following convergence parameters: a maximal bond dimension of $400$ states, which ensured a truncation error of at most $5\times 10^{-8}$ at the final time, a time-step of $dtJ/\hbar=2.5\times 10^{-3}$, the local Hilbert space of the bosonic atoms is $N_{\text{bos}}=4$, and the adaptive cutoff of the local Hilbert space of the photonic mode ranged between $N_{\text{pho}}=70$ and $N_{\text{pho}}=8$. For the parameters used in the presence of dissipation a truncation error of $10^{-8}$ at the final time was achieved for the mentioned convergence parameters, the presented results are averaged over at least 750 quantum trajectories.
The numerical results for the time-evolution of the atomic model with global range interactions, Eq. (4) in the main text, were obtained with an implementation based on matrix product states of the two-site version of the time-dependent variational principle approach (TDVP) Haegeman et al. (2011, 2016), within the ITensor Library Fishman et al. (2022). The convergence was assured by the following parameters of the method: a maximal bond dimension of $250$ states, which ensured a truncation error of at most $10^{-7}$ at the final time, a time-step of $dtJ/\hbar=5\times 10^{-3}$, and the local Hilbert space of the bosonic atoms are $N_{\text{bos}}=4$.
B Approximate Analytical calculations
1 In the limit of small coupling
In this section, we approximate the time-evolution operator in the limit of a small coupling $\hbar\Omega\ll J,U$ and short times $t$. The time-evolution operator is given by $\hat{U}(t)=e^{-i\hat{H}t/\hbar}$. Thus, for the Hamiltonian
$\hat{H}=\hat{H}_{\text{c}}+\hat{H}_{\text{ac}}+\hat{H}_{\text{int}}+\hat{H}_{\text{kin}}$ (B.1)
[…]
$\hat{H}_{\text{c}}=\hbar\delta\hat{a}^{\dagger}\hat{a},~{}\hat{H}_{\text{ac}}=-\hbar\Omega(\hat{a}+\hat{a}^{\dagger})\hat{\Delta},$
by employing the Zassenhaus formula we can rewrite the exponential of the sum of different terms as
$\hat{U}(t)^{\dagger}$ $=e^{i\hat{H}t/\hbar}$ (B.2)
$=e^{i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}e^{i(\hat{H}_{\text{ac}}+\hat{H}_{\text{c}})t/\hbar}$
×$\qquad\times e^{-t^{2}/2\hbar[\hat{H}_{\text{kin}}+\hat{H}_{\text{int}},\hat{H}_{\text{ac}}]}\dots,$
where the dots stand for exponentials of higher order commutators. By expanding the exponentials in the orders of $\Omega$ we obtain
$\hat{U}(t)\approx e^{-i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}+O(t\Omega).$ (B.3)
The term linear in $t\Omega$ vanishes in the evolution of the density-density correlations, since the coupling term $\hat{H}_{ac}$ commutes with the density-density operator, i.e.
$\frac{\partial}{\partial t}$ $\langle\hat{n}_{j}\hat{n}_{j+d}\rangle\propto$ (B.4)
$~{}\big{\langle}e^{i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}\left[1-it\Omega(\hat{a}+\hat{a}^{\dagger})\hat{\Delta}\right]\hat{n}_{j}\hat{n}_{j+d}$
×$~{}~{}\times\left[1+it\Omega(\hat{a}+\hat{a}^{\dagger})\hat{\Delta}\right]e^{-i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}\big{\rangle}+O(t^{2}\Omega)$
$=$ $\big{\langle}e^{i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}\hat{n}_{j}\hat{n}_{j+d}e^{-i(\hat{H}_{\text{kin}}+\hat{H}_{\text{int}})t/\hbar}\big{\rangle}+O(t^{2}\Omega)$
A similar argument holds for the evolution of the average value of the local density. If one takes these expression together, it implies that the time-evolution of the density-density correlations are dominated by the short range terms of the Hamiltonian and obtain only corrections on the order of $O(\Omega t^{2})$ from the global-range couplings. For many different initial states, one can further show that only higher order correction are important in the short-time dynamics, as we show in the following section.
2 Simplified model for the description of the short-time density-density correlation growth
In this section, we aim to capture the minimal ingredients necessary for understanding the distance-independent spreading of the correlations and derive the short-time scaling for the simplified model, $\mathcal{C}_{nn}(d,t)\propto J^{4}\Omega^{4}t^{8}$. As discussed in the main text, when starting from an uncorrelated Fock state the kinetic processes are responsible for creating density-density correlations locally and the coupling to the cavity determines the supersonic spreading. Thus, as a first approximation we neglect the on-site repulsive interactions and the detuning $\delta$ to explain the short time behaviour. In this situation, the 1D version of the Hamiltonian in Eq. (2) of the main text can be written in the momentum basis as
$\hat{H}=-2J\sum_{k\in\mathbb{B}}\cos(k)\hat{b}^{\dagger}_{k}\hat{b}_{k}-\hbar\Omega(\hat{a}+\hat{a}^{\dagger})\hat{\Delta},$ (B.5)
where we assumed periodic boundary conditions and with
$\hat{\Delta}=$ $\sum_{k\in\mathbb{B}}\hat{b}_{k}^{\dagger}\hat{b}_{k+\pi},$ (B.6)
$\hat{b}_{k}=$ $\frac{1}{\sqrt{L}}\sum_{l}e^{ikl}\hat{b}_{l},$ (B.7)
and $\mathbb{B}=\{-\pi,-\pi+2\pi/L,\dots,\pi-2\pi/L\}$, where $L$ is the number of lattice sites.
This Hamiltonian can be rewritten in the form
$\hat{H}=\sum_{k\in\mathbb{B}^{\prime}}\begin{pmatrix}\hat{b}^{\dagger}_{k+\pi}&\hat{b}^{\dagger}_{k}\end{pmatrix}\begin{pmatrix}2J\cos(k)&-\hbar\hat{\Omega}\\ -\hbar\hat{\Omega}&-2J\cos(k)\end{pmatrix}\begin{pmatrix}\hat{b}_{k+\pi}\\ \hat{b}_{k}\end{pmatrix}$ (B.8)
with $\mathbb{B}^{\prime}=\{-\pi/2+2\pi/L,-\pi/2+4\pi/L,\cdots,\pi/2\}$ and
$\hat{\Omega}=$ $\Omega(\hat{a}^{\dagger}+\hat{a}).$ (B.9)
In the following we make an additional assumption by neglecting the full momentum dependence of the kinetic term and replacing it with the constant $J$ in the diagonal entries of Eq. (B.8). This is justified in the regime in which the energy scale $\hbar\Omega$ is much larger than $J$. The resulting Hamiltonian reads
$\hat{H}\approx\sum_{k\in\mathbb{B}^{\prime}}\begin{pmatrix}\hat{b}^{\dagger}_{k+\pi}&\hat{b}^{\dagger}_{k}\end{pmatrix}\begin{pmatrix}J&-\hbar\hat{\Omega}\\ -\hbar\hat{\Omega}&-J\end{pmatrix}\begin{pmatrix}\hat{b}_{k+\pi}\\ \hat{b}_{k}\end{pmatrix}.$ (B.10)
This simplified Hamiltonian can be diagonalized
$H=\sum_{k}\hat{E}[\hat{e}_{k,+}^{\dagger}\hat{e}_{k,+}-\hat{e}_{k,-}^{\dagger}\hat{e}_{k,-}],$ (B.11)
with $\hat{E}=\sqrt{J^{2}+\hbar^{2}\hat{\Omega}^{2}}$ and
$\hat{e}_{k,+}=\cos(\hat{\theta})\hat{b}_{k+\pi}-\sin(\hat{\theta})\hat{b}_{k},$ (B.12)
$\hat{e}_{k,-}=\sin(\hat{\theta})\hat{b}_{k+\pi}+\cos(\hat{\theta})\hat{b}_{k},$ (B.13)
where
$\cos(\hat{\theta})=$ $\frac{J+\hat{E}}{\sqrt{2\hat{E}(\hat{E}+J)}},$ (B.14)
$\sin(\hat{\theta})=$ $\frac{\hat{\Omega}}{\sqrt{2\hat{E}(\hat{E}+J)}}.$ (B.15)
We use this result to derive the time-evolution of the bosonic annihilation and creation operators in real space. In the case of $l$ is even we obtain
$\hat{b}_{l}^{\dagger}(t)=$ $\frac{1}{\sqrt{L}}\sum_{k\in\mathbb{B}^{\prime}}e^{-ikl}[\hat{b}_{k+\pi}^{\dagger}(t)+\hat{b}_{k}^{\dagger}(t)]$
$=$ $\frac{1}{\sqrt{L}}\sum_{k\in\mathbb{B}^{\prime}}e^{-ikl}\cos(\hat{\theta})[e^{i\hat{E}t/\hbar}\hat{e}_{k,+}^{\dagger}+\hat{e}_{k,-}^{\dagger}e^{-i\hat{E}t/\hbar}]$
$-\frac{1}{\sqrt{L}}\sum_{k\in\mathbb{B}^{\prime}}e^{-ikl}\sin(\hat{\theta})[e^{i\hat{E}t/\hbar}\hat{e}_{k,+}^{\dagger}-\hat{e}_{k,-}^{\dagger}e^{-i\hat{E}t/\hbar}]$
$=$ $\left[\cos(\hat{E}t/\hbar)-i\frac{\hbar\hat{\Omega}\sin(\hat{E}t/\hbar)}{\hat{E}}\right]\hat{b}_{l}^{\dagger}(0),$ (B.16)
where we used Eqs. (B.12) and (B.13) and ${\sin(2\hat{\theta})=\hbar\hat{\Omega}/\hat{E}}$. Consequently, we can calculate for two even sites $l,m$ the density-density correlations with
$\mathcal{C}_{nn}(|l-m|,t)=$ $\left\langle\hat{n}_{l}\hat{n}_{m}\right\rangle-\left\langle\hat{n}_{l}\right\rangle\left\langle\hat{n}_{m}\right\rangle$
$=$ $\mathrm{Var}(\hat{C})\langle\hat{n}_{l}(0)\rangle\langle\hat{n}_{m}(0)\rangle,$ (B.17)
with
$\hat{C}=$ $\cos^{2}(\hat{E}t/\hbar)+\frac{\hbar^{2}\hat{\Omega}^{2}\sin^{2}(\hat{E}t/\hbar)}{\hat{E}^{2}},$ (B.18)
$\mathrm{Var}(\hat{C})=$ $\langle\hat{C}^{2}\rangle-\langle\hat{C}\rangle^{2}.$ (B.19)
Here we used that the initial state is a product of the photonic state and a Fock state of the atoms in order to decouple the expectation values in Eq. (2).
The final result now follows from the Taylor expansion of $\hat{C}$ for small times
$\hat{C}\approx 1-J^{2}t^{2}/\hbar^{2}+\frac{J^{4}t^{4}}{3\hbar^{4}}+\frac{J^{2}\hat{\Omega}^{2}t^{4}}{3\hbar^{2}}$ (B.20)
and therefore
$\mathrm{Var}(\hat{C})$ $\approx\frac{J^{4}\Omega^{4}t^{8}}{9\hbar^{2}}\mathrm{Var}([\hat{a}+\hat{a}^{\dagger}]^{2}),$ (B.21)
This equation predicts the growth of density-density correlations ${\mathcal{C}_{nn}\propto J^{4}\Omega^{4}t^{8}}$ and shows the importance of the fluctuations of the cavity field operators.
C Treating the cavity mode as a classical field
Figure C1: (a), (c), (e) The space-time propagation of the correlations $\mathcal{C}_{nn}(d,t)$, in the presence of dissipation, and (b), (d), (f) the time-dependence of $\mathcal{C}_{nn}(d,t)$ for several distances. In panels (a), (b) we show the full quantum dynamics of the atoms-cavity model, Eqs. (1)-(2) in the main text, in panels (c), (d) the cavity dynamics has been replaced with a classical field, Eq. (C.1) and Eq. (C.2), and in panels (e), (f) stochastic noise has been added to the classical field evolution, Eq. (C.1) and Eq. (C.3). The blue lines are a guide to the eye and approximates the front of the light-cone propagation for $\Omega=0$ and $\Gamma=0$. The magenta lines represent algebraically increasing curves $\propto t^{\alpha}$, with $\alpha\in\{2,4,8\}$. The parameters used are $\hbar\Omega/J=0.99$ $\hbar\Gamma/J=10$, $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$. Panels (a) and (c) correspond to Fig. 3(a) and Fig. 3(b) in the main text, we reproduce them here for completeness.
In the main text we contrast the propagation dynamics of the correlations for the Bose-Hubbard model coupled to the quantum dissipative field of an optical cavity with results in which the cavity mode has been replaced with a classical field. The classical field realizes a superlattice staggered potential.
The classical staggered potential can be derived as a mean-field description of the cavity-atoms coupling approach in which the cavity field is described by a classical coherent field Ritsch et al. (2013); Maschler et al. (2008); Nagy et al. (2008). For the one-dimensional case, within this approximation, the atoms are described by the Hamiltonian
$\hat{H}_{\text{MF}}=\hat{H}_{\text{int}}+\hat{H}_{\text{kin}}+\hat{H}_{\text{stag}}$ (C.1)
$\hat{H}_{\text{int}}=\frac{U}{2}\sum_{j=1}^{L}\hat{n}_{j}(\hat{n}_{j}-1),$
$\hat{H}_{\text{kin}}=-J\sum_{j=1}^{L-1}(\hat{b}_{j}^{\dagger}\hat{b}_{j+1}+\hat{b}_{j+1}^{\dagger}\hat{b}_{j}),$
$\hat{H}_{\text{stag}}=-V(t)\hat{\Delta},~{}~{}\hat{\Delta}=\sum_{j=1}^{L}(-1)^{j}\hat{n}_{j}.$
Here $V(t)$ is a time-dependent field, describing the coupling to cavity mode which is assumed to be in a coherent state, $V(t)=\Omega\left\langle\hat{a}^{\dagger}+\hat{a}\right\rangle(t)$, with the equation of motion
$\frac{\partial}{\partial t}\langle\hat{a}\rangle=i\Omega\langle\hat{\Delta}\rangle-\left(i\delta+\Gamma/2\right)\langle\hat{a}\rangle.$ (C.2)
We integrate this equation of motion in a coupled way with the time-evolution of $H_{\text{MF}}$, where the potential coupled to the atomic odd-even imbalance, $V(t)$, depends on the imbalance $\langle\hat{\Delta}\rangle(t-\text{d}t)$ at the previous time step.
The comparison between the exact results and the classical field approach is also shown in Fig. 3. As discussed in the main text, by the coupling to the cavity field for the parameters shown the dynamics of the correlations is dominated by a supersonic propagation [Fig. 3(a) and Fig. 3(b)], while for the classical field approach we only observe a light-cone spreading of the correlations [Fig. 3(c) and Fig. 3(d)].
In order to understand the necessary ingredients for the supersonic propagation of the correlations, we also contract our results with the classical field approach for describing the cavity field to which stochastic noise has been added to its dynamics
$\frac{\partial}{\partial t}\langle\hat{a}\rangle=i\Omega\langle\hat{\Delta}\rangle-\left(i\delta+\Gamma/2\right)\langle\hat{a}\rangle+\sqrt{\Gamma}\xi(t),$ (C.3)
where $\xi(t)$ is a random complex number of magnitude $1$ sampled from a uniform distribution at each point in time. For the initial state $\langle a\rangle(0)$ we use a uniformly sampled random complex number. We simulate different realizations of the time-dependent noise term, similar to the quantum trajectories approach, and average over the values of the computed observables for the different realizations. We obtain that in the presence of the stochastic noise, Fig. C1(e) and Fig. C1(f) it exhibits for larger distances features of the distance-independent rise of the correlations, as seen in Fig. C1(f). This can be understood by the fact that the noise term appearing in Eq. (C.3) translates to a correlated noise term in the evolution of the atoms acting on all sites. However, the contribution from the distance-independent rise of the correlations is much smaller for the shown parameters compared to the full quantum dynamics. This implies that the noise term is insufficient to capture the full dynamics which requires, in particular, also the cavity-mediated transport of density fluctuations.
D Truncated Wigner approximation
In this section, we describe the truncated Wigner simulation methods Gardiner et al. (2002); Olsen et al. (2004); Nagao et al. (2019) and we provide additional information regarding the truncated Wigner results presented in the main texts. The truncated Wigner simulation is based on semiclassical equations of motion for the bosonic field operators describing the atomic and the cavity degrees of freedom. For each lattice site $j$ we have two real fields $\beta^{(r)}_{j}$ and $\beta^{(i)}_{j}$ that can be captured as one complex field $\beta_{j}=\beta^{(r)}_{j}+i\beta^{(i)}_{j}$ representing the bosonic operator $\hat{b}_{j}$. In addition, we have two real cavity fields $\alpha^{(r)}$ and $\alpha^{(i)}$ that define a complex cavity field $\alpha=\alpha^{(r)}+i\alpha^{(i)}$ representing the bosonic cavity field operator $\hat{a}$.
In order to derive the semiclassical equations of motions we start with the Heisenberg-Langevin equations of motion for the operators that are given by
$\frac{d\hat{b}_{j}}{dt}=$ $i\frac{J}{\hbar}\sum_{l\in\langle l,j\rangle}\hat{b}_{l}-i\frac{U}{\hbar}\hat{b}_{j}^{\dagger}\hat{b}_{j}\hat{b}_{j}$
$+i\Omega(\hat{a}+\hat{a}^{\dagger})(-1)^{j}\hat{b}_{j}$ (D.1)
$\frac{d\hat{a}}{dt}=$ $-\left(i\delta+\frac{\Gamma}{2}\right)\hat{a}+\sqrt{\Gamma}\hat{a}_{\mathrm{in}}(t)$ (D.2)
$+i\Omega\sum_{j}(-1)^{j}\hat{n}_{j},$
where $\langle l,j\rangle$ is the set of neighboring sites of $j$. The input shot noise $\hat{a}_{\mathrm{in}}(t)$ has vanishing mean value $\langle\hat{a}_{\mathrm{in}}(t)\rangle$ and second moments $\langle\hat{a}_{\mathrm{in}}(t)\hat{a}_{\mathrm{in}}(t^{\prime})\rangle=0=\langle\hat{a}^{\dagger}_{\mathrm{in}}(t)\hat{a}_{\mathrm{in}}(t^{\prime})\rangle$, $\langle\hat{a}_{\mathrm{in}}(t)\hat{a}^{\dagger}_{\mathrm{in}}(t^{\prime})\rangle=\delta(t-t^{\prime})$.
To obtain the semiclassical equations of motion for the real fields $\alpha^{(r)}$, $\alpha^{(i)}$ and $\beta^{(r)}_{j}$, $\beta^{(i)}_{j}$, we first derive the equations of motion for $\hat{b}_{j}^{(r)}=(\hat{b}_{j}+\hat{b}_{j}^{\dagger})/2$, $\hat{b}_{j}^{(i)}=(\hat{b}_{j}-\hat{b}_{j}^{\dagger})/(2i)$ and $\hat{a}^{(r)}=(\hat{a}+\hat{a}^{\dagger})/2$, $\hat{a}^{(i)}=(\hat{a}-\hat{a}^{\dagger})/(2i)$. Subsequently we perform a symmetric ordering of the operators. Note that for instance $\hat{n}_{j}=\hat{b}_{j}^{\dagger}\hat{b}_{j}=[\hat{b}_{j}^{(r)}]^{2}+[\hat{b}_{j}^{(i)}]^{2}-1/2$. After that we exchange the operators $\hat{b}_{j}^{(r)}$, $\hat{b}_{j}^{(i)}$ by the real variables $\beta^{(r)}_{j}$, $\beta^{(i)}_{j}$ and $\hat{a}^{(r)}$, $\hat{a}^{(i)}$ by the real variables $\alpha^{(r)}$, $\alpha^{(i)}$. With these equations of motion we can write down a set of complex coupled stochastic differential equations for $\alpha=\alpha^{(r)}+i\alpha^{(i)}$ and $\beta_{j}=\beta^{(r)}_{j}+i\beta^{(i)}_{j}$ given by
$\frac{d\beta_{j}}{dt}=$ $i\frac{J}{\hbar}\sum_{l\in n(j)}\beta_{l}-i\frac{U}{\hbar}\left(|\beta_{j}|^{2}-\frac{1}{2}\right)\beta_{j}$ (D.3)
$+i\Omega(\alpha+\alpha^{*})(-1)^{j}\beta_{j}$
$\frac{d\alpha}{dt}=$ $-\left(i\delta+\frac{\Gamma}{2}\right)\alpha+\sqrt{\Gamma}\mathcal{F}(t)$ (D.4)
$+i\Omega\sum_{j}(-1)^{j}\left(|\beta_{j}|^{2}-\frac{1}{2}\right).$
with $\mathcal{F}(t)=[\mathcal{F}_{r}(t)+i\mathcal{F}_{i}(t)]/2$, $\langle\mathcal{F}_{r}\rangle=0=\langle\mathcal{F}_{i}\rangle$, $\langle\mathcal{F}_{r}(t^{\prime})\mathcal{F}_{i}(t)\rangle=0$, and $\langle\mathcal{F}_{i}(t^{\prime})\mathcal{F}_{i}(t)\rangle=\delta(t-t^{\prime})$, $\langle\mathcal{F}_{r}(t^{\prime})\mathcal{F}_{r}(t)\rangle=\delta(t-t^{\prime})$. Equations (D.3) and (D.4) are the dynamical equations we simulate in our truncated Wigner approach with a given initial condition.
1 Initial conditions
To model the initial condition of the atomic Fock state and the empty cavity states used in the simulations of the main text, we initialize the cavity in the vacuum state and sample $\alpha^{(r)}(t=0)$ and $\alpha^{(i)}(t=0)$ with $\alpha(t=0)=\alpha^{(r)}(0)+i\alpha^{(i)}(0)$ as Gaussian random variables with $\langle\alpha^{(r)}(0)\rangle=0=\langle\alpha^{(i)}(0)\rangle$, $\langle\alpha^{(r)}\alpha^{(i)}\rangle=0$, and $\langle\alpha^{(r)}(0)\alpha^{(r)}(0)\rangle=1/4=\langle\alpha^{(i)}(0)\alpha^{(i)}(0)\rangle$. For the atomic state we assume that each mode indexed by $j$ is in a Fock state with occupation $n_{j}$. For this we employ a sampling method which was previously derived in Refs. Gardiner et al. (2002); Olsen et al. (2004); Nagao et al. (2019). We sample $\beta_{j}(0)=[p_{j}+q_{j}\eta_{j}]e^{i\varphi_{j}}$ where $\eta_{j}$ is a Gaussian random variable with mean equal zero and with variance $\langle\eta_{j}^{2}\rangle=1$ and $\varphi_{j}$ is a uniform random variable in the interval $[0,2\pi)$. The values of $p_{j}$ and $q_{j}$ are determined by the occupation $n_{j}$ with
$p_{j}=$ $\frac{\sqrt{2n_{j}+1+2\sqrt{n_{j}^{2}+n_{j}}}}{2},$ (D.5)
$q_{j}=$ $\frac{1}{4p_{j}}.$ (D.6)
Note that with this definition we find
$\langle\beta_{j}^{*}(0)\beta_{j}(0)\rangle=$ $n_{j}+\frac{1}{2}$ (D.7)
and
$\langle\beta_{j}^{*}(0)\beta_{j}^{*}(0)\beta_{j}(0)\beta_{j}(0)\rangle=$ $\left(n_{j}+\frac{1}{2}\right)^{2}+\frac{1}{4}.$ (D.8)
2 Technical details on the simulations
For the simulation of the stochastic differential equation we have used a Runge-Kutta method of 4th order with an integration time step $dt=0.5\times 10^{-2}/\Gamma$. All results shown in the main text are obtained from $M=10^{6}$ noise initializations. For each of the initialization that we index in the following by $m=1,...,M$ we obtain fluctuating densities
$n^{(m)}_{j}(t)=|\beta^{(m)}_{j}(t)|^{2}-\frac{1}{2}.$ (D.9)
The approach is then slighty different for the 1D and 2D cases.
1D simulation:
For the 1D simulation we choose $j=1,\dots,L$ with $L=21$ and we calculate the density-density correlations with
$\mathcal{C}_{nn}(d,t)=$ $\frac{1}{M}\sum_{m=1}^{M}n^{(m)}_{j_{0}}(t)\frac{n^{(m)}_{j_{0}+d}(t)+n^{(m)}_{j_{0}-d}(t)}{2}$ (D.10)
$-\frac{1}{M^{2}}\sum_{m,m^{\prime}=1}^{M}n^{(m^{\prime})}_{j_{0}}(t)\frac{n^{(m)}_{j_{0}+d}(t)+n^{(m)}_{j_{0}-d}(t)}{2}$ (D.11)
where $j_{0}=(L+1)/2$ is the center and the division by $2$ comes from the fact that there are two sites with the same distance in 1D. The result of this calculation is shown in Fig. 3(c) in the main text.
2D simulation:
For the 2D simulation we use ${\bf j}=(j_{x},j_{y})$ with $j_{x}=1,\dots,L$ and $j_{y}=1,\dots,L$ and $L=21$. The center is then $\mathbf{j}_{0}=([L+1]/2,[L+1]/2)$.
We can then calculate the density-density correlations with
$\mathcal{C}_{nn}(d,t)=$ $\frac{1}{M}\sum_{m=1}^{M}n^{(m)}_{\mathbf{j}_{0}}(t)\sum_{{\bf j}}\frac{n^{(m)}_{{\bf j}}(t)}{\mathcal{N}_{d}}\delta_{d({\bf j}),d}$ (D.12)
$-\frac{1}{M^{2}}\sum_{m^{\prime}=1}^{M}n^{(m^{\prime})}_{\mathbf{j}_{0}}(t)\sum_{m=1}^{M}\sum_{{\bf j}}\delta_{d({\bf j}),d}\frac{n^{(m)}_{{\bf j}}(t)}{\mathcal{N}_{d}},$
where $d({\bf j})=\sqrt{(j_{x}-[L+1]/2)^{2}+(j_{y}-[L+1]/2)^{2}}$ is the distance from the center and $\mathcal{N}_{d}$ is the number of sites with distance $d({\bf j})=d$. The result of this calculation is shown in Fig. 3(d) of the main text.